Show that the maximum likelihood estimator of the parameter theta of a population having the density function (2/ theta^2) (theta - x), 0 < x < theta, for a sample of unit size is 2x, where x is the sample value. Also, show that the estimate obtained is biased.
1
Expert's answer
2020-09-08T15:08:20-0400

"f(x,\theta)=\frac{2(\theta-x)}{\theta^2}" , "0<x<\theta"

"L(x,\theta)=\frac{2(\theta-x)}{\theta^2}"

"\frac{\partial L(x,\theta)}{\partial \theta}=\frac{-2}{\theta^2}(1-\frac{2x}{\theta})"

"\frac{-2}{\theta^2}(1-\frac{2x}{\theta})=0"

"1=\frac{2x}{\theta}"

"2x=\hat\theta" Thus, the MLE of "\theta" is 2x.

To check for biasedness, expectation of 2x shoud not be equal to theta.

"E(2x)=\smallint_0^{\theta}\frac{2x(2(\theta-x))}{\theta^2}dx"

"=[\frac{2x^2(3\theta-2x)}{3\theta^2}]_0^\theta"

"=\frac{2\theta^2(3\theta-2\theta)}{3\theta^2}-0"

"=\frac{2\theta}{3}\ne\theta" , Thus, the MLE (2x) is a biased estimator of theta.